# Quantum Data Centers: Why Entanglement Changes Everything

### v1

@@ 3.1.1 Direct vs Entanglement-Generation Transduction
For an in-depth treatise, we refer the reader to [31]. The same hardware to implement a quantum transducer can be exploited for performing two distinct functionalities:
@@
Different transducer-based archetypes, which exploit the different functionalities of quantum transduction, can be envisioned for interconnecting superconducting processors, as schematically summarized in Fig. 4. It is worthwhile to note that when DQT acts on informational qubits, it enables the direct transmission of quantum information. However, Section 1.1 highlighted the fundamental limitations of direct transmission in distributed quantum computation architectures, arising from the no-cloning theorem and quantum measurement postulate [8]. Indeed, the challenges in DQT on informational qubits are even more critical than standard direct quantum transmission, since transduction errors compound with transmission losses. In fact, the quantum transduction should be able to preserve the encoded quantum information in the conversion, that constitutes a not trivial task. Consequently, distributed quantum computing architectures must rely on entanglement-mediated approaches, such as DQT on ebits or EGT.
@@
DQT on ebits is exploited for entanglement distribution. Specifically, the entanglement resource is locally generated and then distributed to the remote node via quantum transduction. Once the entanglement is successfully generated and distributed, remote operations can be executed accordingly to the TeleData and TeleGate paradigms as in Fig. 4(b). The main advantage of DQT on ebits lies in the possibility of entanglement re-generation [27]. Indeed, if the ebits are lost in the conversions (up- at source or down- at destination), they can be regenerated without restrictions, being a communication resource rather than an information. The re-generation process can be performed until the entanglement is successfully distributed between remote nodes [31].
@@
In the EGT paradigm, the transducer generates hybrid entanglement between different frequencies domains [31, 33, 45, 46, 47], i.e., entanglement between a microwave photon and an optical photon in the form of a 2-level Fock state [^1] The approximation of a two-level Fock-state is satisfied with some hypothesis regarding the physical interaction underlying transduction. We refer the interested reader to [31] and [34].:
|$\ket{\Psi_{m,o}}=\alpha\ket{0_{m}1_{o}}+\beta\ket{1_{m}0_{o}},$ (1)
@@
where the subscripts $m$ and $o$ refer to the frequency domain of each ebit, namely microwave and optical, and $\alpha$ and $\beta$ depends on the transducer hardware [34, 30]. EGT can be exploited to implement the Duan–Lukin–Cirac–Zoller protocol [48], represented in Fig. 4(c). Accordingly, the optical ebits generated in each node with EGT travel through the optical fibers to reach an optical repeater node. The repeater is equipped with a 50/50 beam splitter followed by two detectors. The click of one detector denotes the presence of an optical photon. In this configuration when a detector clicks, it is impossible to identify from which transducer the optical photon comes from, or, equivalently, it is impossible to distinguish whether a microwave photon is present at source or at destination. Therefore, the overall effect of beam splitter and detectors is reminiscent of entanglement swapping, projecting the received optical photons into a Bell state [31, 30]. This results into the generation of path-entanglement [49] between the microwave photons at the source and at the destination:
|$\ket{\Psi_{m,m}^{s,d}}=\frac{1}{\sqrt{2}}(\ket{0_{m}^{s}1_{m}^{d}}+\ket{1_{m}^{s}0_{m}^{d}}),$ (2)
@@
where the up-scrips $s$ and $d$ denote the “location" of each ebit, namely at source and destination. Therefore, this archetype for entanglement distribution exploiting two EGTs and entanglement swapping allows to herald entanglement between remote superconducting devices, without destroying it. Once the distributed entanglement is heralded, it can be leveraged for remote operations.
@@
The key advantage of EGT over DQT on ebits is to relax the constrains on the conversion efficiency $\eta$ – the main hardware parameter characterizing the transducer performance – for having a non-null entanglement distribution probability, as depicted in Fig. 5.
@@
The described framework can be adopted in a Quantum Data Center scenario, where – as aforementioned – a specialized node, the orchestrator, is responsible for entanglement generation and distribution to the other nodes [34, 30]. This generalization can also include in the picture the possibility of generating not only bipartite entangled states but also multipartite entangled states, allowing new and astonishing network functionalities. This has been deeply discussed from a quantum transduction perspective in [34, 30].
@@ 3.2 QLAN architecture: physical vs artificial topology
As aforementioned, in Quantum Data Center archetype, the quantum processors are interconnected via a quantum local area network (QLAN). We clarified at the beginning of Section 3 that the quantum technology maturity imposes on one hand sparse physical QLAN topologies and, on the other hand, a hierarchical QLAN design where specialized nodes, the orchestrator nodes, are responsible for the entanglement generation and distribution. Consequently, the network logic is centralized at the orchestrator node, while the remaining nodes, referred in the following as client nodes, are lightweight and simple [23, 22]. Although this design fundamentally departures from classical network paradigms, the quantum-specific constraints, including decoherence timescales and hardware complexity, make this centralized approach viable.
Figure 6: Physical star topology of a QLAN with $6$ client nodes and a single orchestrator (left). Examples of different artificial topologies obtainable by proper manipulations of the shared multipartite entangled state (right).
@@
The sparse QLAN physical connectivity inherently restricts the communication capabilities between the QLAN nodes. For overcoming these limitations, we cannot borrow well-established approaches from classical LANs [22], where topological constraints are overcome through the upper layers of the protocol stack, at the price of communication overhead and information duplication. This impossibility is due to the unconventional quantum peculiarities, ranging from stringent coherence times to quantum mechanics postulates and phenomena, such as quantum measurement and the no-cloning theorem. Furthermore, the design of a protocol suite for quantum networks is still at its infancy [18], and thus the functionalities of “quantum” upper-layers are yet to be defined.
@@
Fortunately, the communication limitations induced by the physical QLAN topology can be overcome, by relying on quantum entanglement, which enables a fundamentally new form of connectivity, referred to as entanglement-based connectivity [27, 23, 22]. Specifically, by leveraging multipartite entanglement, it is possible to establish a virtual topology – referred to as artificial topology – overlaying the physical one. The artificial topologies activated by entanglement can significantly differ from the underlying physical topology, and they can – if properly engineered – effectively mitigate the constraints imposed by the QLAN architecture.
@@
Indeed, as depicted in Fig. 6, the required artificial topology can be engineered to support a wide range of quantum communication patterns. However, the choice and the design of the multipartite entangled state to be generated and distributed is highly non-trivial and has great influence on the enabled network functionalities [29, 29, 28]. As instance, it is well known that by sharing a GHZ state, a single Bell state can be extracted, on-demand, whenever needed, between any pair of network nodes sharing part of the entangled state [29]. This has the advantage to dynamically adapt to the communication demands, without predetermining the identities of the involved nodes [^2] Conversely, when a Bell state is shared between two parties, the identities of nodes involved in the communication has been fixed a-priori.. Although GHZ states present interesting properties compared to the use of bipartite entangled states, they are limited by the extraction of a single Bell state, even if the number of qubits of the overall state grows arbitrarily. At the same time, GHZ states are characterized by unitary persistency [50], i.e., the minimum number of qubits that need to be measured to guarantee that the resulting state is separable is one. Thus, they are highly sensitive to errors: losses of even one of the particles causes the entire state to lose coherence.
@@
Accordingly, the choice of the multipartite state to be distributed in the network involves fundamental compromises between three main factors: the number of Bell pairs that can be extracted – after manipulations of the initial state – persistency and robustness to quantum noise. Thus the degrees of freedom for the design can be resumed as follows [22]:
@@
Pauli measurements on |$\ket{G}$ Corresponding graph operations
$P_{z,\pm}^{(a)}\ket{G}=\ket{z,\pm}^{(a)}\otimes\underbrace{U_{z,\pm}^{(a)}\ket{G-a}}_{\ket{\tilde{G}_{z}}}$ $\tilde{G}_{z}=G-a$
$P_{y,\pm}^{(a)}\ket{G}=\ket{y,\pm}^{(a)}\otimes\underbrace{U_{y,\pm}^{(a)}\ket{\tau_{a}(G)-a}}_{\ket{\tilde{G}_{y}}}$ $\tilde{G}_{y}=\tau_{a}(G)-a$
$P_{x,\pm}^{(a)}\ket{G}=\ket{x,\pm}^{(a)}\otimes\underbrace{U_{x,\pm}^{(a)}\ket{\tau_{b_{0}}\left(\tau_{a}\left(\tau_{b_{0}}(G)\right)-a\right)}}_{\ket{\tilde{G}_{x}}}$ $\tilde{G}_{x}=\tau_{b_{0}}\big{(}\tau_{a}(\tau_{b_{0}}(G))-a\big{)}$
Table 2: Table of single-qubit Pauli measurements with projectors $P_{i,\pm}^{(a)},\,i\in\{x,y,z\}$ acting on a generic qubit $a$ of the graph state |$\ket{G}$ (with outcome ±$\pm 1$) and corresponding graph operations, in terms of vertex deletions and local complementations $\tau(\cdot)$. Table reproduced from [24].
@@ 3.2.1 Graph states as network resources
Stemming from the above discussion, graph states are particularly interesting candidates, due to their unique entanglement properties, noise robustness and intrinsic resilience to particle losses [51, 52, 53, 32, 54]. Interestingly, graph states have a straightforward representation of entanglement interactions in the form of a graph. In other words, each graph state |$\ket{G}$ has its own associated graph $G=(V,E)$, and can be written in the following form:
|$\ket{G}=\prod_{\{a,b\}\in E}\texttt{CZ}_{ab}\ket{+}^{\otimes n},$ (3)
where |$\ket{+}=\tfrac{1}{\sqrt{2}}(\ket{0}+\ket{1})$ and where the controlled-Z gate ($\mathtt{CZ}$) represents the entangling operation between qubits $a$ and $b$, corresponding to an edge in the associated graph $G=(V,E)$. One of the most interesting properties of graph states is the possibility to engineer the state structure, by wisely applying local quantum operations on some qubits. More into details, the application of single-qubit Pauli measurements – denoted by the projection operator $P_{i}$ with $i\in\{x,y,z\}$ – map into specific graph operations on the associated graph[51, 52]. As explicitly reported in Tab 2 and visually represented in Fig. 7(a), the graph state obtained after each Pauli measurement is equivalent – up to the application of unitary operators $U_{i,\pm}$ – to a new graph state, whose associated graph is obtained through vertex deletions and local complementations $\tau(\cdot)$ on the original graph $G$ [52, 22, 24].
The simplest graph state is the $n$-qubits linear graph state, whose associated graph is a linear interconnection of vertices and its expression reduces to: |$\ket{L}=\prod_{i=1}^{n-1}\texttt{CZ}_{i,i+1}\ket{+}^{\otimes n}$. Such a simple graph state is particularly interesting in the context of quantum communication and computation [55, 56], since it can be used as building block for more complex target states [57], according to proper manipulations and merging of multiple instances, as experimentally proven in controlled environments [58, 59, 60].
@@
Figure 7: Schematic representation of the correspondence between graph states measurements and the associated graph operations (a). Application of Pauli-Y measurements for the engineering of the artificial topology of the network in a $6$ clients QLAN topology (b).
@@ 3.2.2 From physical to artificial topologies
Accordingly to the above, graph states emerge as promising resource for engineering the QLAN topologies, due to their unique entanglement properties and operational flexibility[23, 24, 22]. The centralized QLAN model, depicted in Fig. 6, allows the orchestrator node to generate and distribute graph states and linear graph states, while minimizing client-side operations and associated communication overhead. To this aim, it retains for itself part of the qubits of the generated graph state while distributing the remaining ones to the client nodes. With this strategy, the orchestrator node can engineer the artificial topology solely through the use of its locally held qubits and by employing Local Operations and Classical Communication (LOCC). Thus, this approach allows the dynamic reconfiguration of the QLAN topology, without requiring distributed quantum operations at the clients, besides the correction unitaries after the Pauli measurements, as indicated in Tab 2.
@@
The orchestrator can leverage the inherent simplicity of linear graph states as building block for synthesizing more sophisticated entangled states, as recently shown in [22]. This approach enables the engineering of denser QLAN artificial topologies. This, in turn, implies that the orchestrator gains enhanced capabilities to tailor the connectivity patterns, allowing for more versatile and resource-efficient quantum computation architectures.
An example of dense resulting artificial topology is represented in Fig. 6(b). Remarkably, the presence of multiple entangled links in the artificial topology does not necessarily imply the extraction of more Bell states [22], but rather simplifies the selection of the identities of the target nodes for dedicated link extractions. As a consequence, the orchestrator node is capable of enhancing the reliability and adaptability of the overlying entangled network topology, thanks to the possibility of reconfiguring paths according to the traffic demands – requests of specific entangled links – or node disconnections. Interestingly, as also represented in Fig. 8(a), the orchestrator is able to generate and distribute wisely manipulated linear graph states through merging [^3] The merging operation between vertices belonging to different graph states combines two vertices into a single one. This operation is equivalent to the so-called fusion when referred to optical graph state operations. operations. Specifically, the orchestrator is able to further engineer the resulting connectivity of the network nodes according to proper Pauli-X measurements. As depicted in Fig. 8(b), the adjacency between any two target clients can be achieved by reducing the distance of the clients in the resulting artificial topology. As a result, proximity engineering of physically unconnected clients has the direct consequence of simplifying the extraction of the dedicated entangled links in the artificial topology.

### v2

@@ 3.1.1 Direct vs Entanglement-Generation Transduction
For an in-depth treatise, we refer the reader to [CalDavHan-25]. The same hardware to implement a quantum transducer can be exploited for performing two distinct functionalities:
@@
Different transducer-based archetypes, which exploit the different functionalities of quantum transduction, can be envisioned for interconnecting superconducting processors, as schematically summarized in Fig. 4. It is worthwhile to note that when DQT acts on informational qubits, it enables the direct transmission of quantum information. However, Section 1.1 highlighted the fundamental limitations of direct transmission in distributed quantum computation architectures, arising from the no-cloning theorem and quantum measurement postulate [CacCalVan-20]. Indeed, the challenges in DQT on informational qubits are even more critical than standard direct quantum transmission, since transduction errors compound with transmission losses. In fact, the quantum transduction should be able to preserve the encoded quantum information in the conversion, that constitutes a not trivial task. Consequently, distributed quantum computing architectures must rely on entanglement-mediated approaches, such as DQT on ebits or EGT.
@@
DQT on ebits is exploited for entanglement distribution. Specifically, the entanglement resource is locally generated and then distributed to the remote node via quantum transduction. Once the entanglement is successfully generated and distributed, remote operations can be executed accordingly to the TeleData and TeleGate paradigms as in Fig. 4(b). The main advantage of DQT on ebits lies in the possibility of entanglement re-generation [IllCalMan-22]. Indeed, if the ebits are lost in the conversions (up- at source or down- at destination), they can be regenerated without restrictions, being a communication resource rather than an information. The re-generation process can be performed until the entanglement is successfully distributed between remote nodes [CalDavHan-25].
@@
In the EGT paradigm, the transducer generates hybrid entanglement between different frequencies domains [CalDavHan-25, DavZhaChu-24, DavCalWan-23, MeeLakWoo-24, SahQiuHea-23], i.e., entanglement between a microwave photon and an optical photon in the form of a 2-level Fock state [^3] The approximation of a two-level Fock-state is satisfied with some hypothesis regarding the physical interaction underlying transduction. We refer the interested reader to [CalDavHan-25] and [DavCacCal-24].:
[…]
@@
where the subscripts $m$ and $o$ refer to the frequency domain of each ebit, namely microwave and optical, and $\alpha$ and $\beta$ depends on the transducer hardware [DavCacCal-24, DavCacCal-2024]. EGT can be exploited to implement the Duan–Lukin–Cirac–Zoller protocol [DuaLukCir-01], represented in Fig. 4(c). Accordingly, the optical ebits generated in each node with EGT travel through the optical fibers to reach an optical repeater node. The repeater is equipped with a 50/50 beam splitter followed by two detectors. The click of one detector denotes the presence of an optical photon. In this configuration when a detector clicks, it is impossible to identify from which transducer the optical photon comes from, or, equivalently, it is impossible to distinguish whether a microwave photon is present at source or at destination. In other words, the overall setup is unable to distinguish the so-called which-path information, that, by oversimplifying, is related to the knowledge about the path a photon takes through a multipath optical system, such as a beam splitter.
This results into the generation of path-entanglement [MontVivCap-15] between the microwave photons at the source and at the destination:
[…]
where the up-scrips $s$ and $d$ denote the “location" of each ebit, namely at source and destination. To summarize, the shared entangled pair in (2) between two microwave photons at the source and destination is obtained from one entangled state between an optical photon and a microwave photon at the source and one entangled state between an optical photon and a microwave photon at the destination. Consequently, the overall result is reminiscent of entanglement swapping[CalDavHan-25, DavCacCal-2024].
@@
This archetype for entanglement distribution exploiting two EGTs and entanglement swapping allows to herald entanglement between remote superconducting devices, without destroying it. Once the distributed entanglement is heralded, it can be leveraged for remote operations.
@@
The key advantage of EGT over DQT on ebits is to relax the constraints on the conversion efficiency $\eta$ – the main hardware parameter characterizing the transducer performance – for having a non-null entanglement distribution probability, as depicted in Fig. 5.
@@
The described framework can be adopted in a Quantum Data Center scenario, where – as aforementioned – a specialized node, the orchestrator, is responsible for entanglement generation and distribution to the other nodes [DavCacCal-24, DavCacCal-2024]. This generalization can also include in the picture the possibility of generating not only bipartite entangled states but also multipartite entangled states, allowing new and astonishing network functionalities. This has been deeply discussed from a quantum transduction perspective in [DavCacCal-24, DavCacCal-2024].
@@ 3.2 QLAN architecture: physical vs artificial topology
As aforementioned, in the Quantum Data Center archetype, the quantum processors are interconnected via a quantum local area network (QLAN). Densely physical-connected QLAN topologies are not practical in a short term. The rationale is that the deployed architectures have to take into account the high cost of data buses (economically and in terms of quantum fidelity) limiting both the size of the clusters and the use of connections for processing. Indeed, as detailed in the previous section, the connections in such controlled environments must either resort to crycables (expensive and necessarily limited in length, for handling the decoherence effects) or resort to quantum transduction, which is still extremely inefficient. Thus, as clarified at the beginning of Section 3, the maturity level of quantum technology leads to two main characteristics of the QLAN. First, the physical topology is sparse. Second, there is a hierarchical structure within the QLAN, where specialized nodes, the orchestrator nodes, are responsible for the entanglement generation and distribution. Consequently, the network logic is centralized at the orchestrator node, while the remaining nodes, referred in the following as client nodes, are lightweight and simple [MazCalCac-24-QCNC, MazCalCac-25]. Although this design fundamentally departures from classical network paradigms, the quantum-specific constraints, including decoherence timescales and hardware complexity, make this centralized approach viable.
Figure 6: Physical star topology of a QLAN with 66 client nodes and a single orchestrator (left). Examples of different artificial topologies obtainable by proper manipulations of the shared multipartite entangled state (right).
@@
A sparse QLAN physical connectivity inherently restricts the communication capabilities between the QLAN nodes. For overcoming these limitations, we cannot borrow well-established approaches from classical LANs [MazCalCac-25], where topological constraints are overcome through the upper layers of the protocol stack, at the price of communication overhead and information duplication. This impossibility is due to the unconventional quantum peculiarities, ranging from stringent coherence times to quantum mechanics postulates and phenomena, such as quantum measurement and the no-cloning theorem. Furthermore, the design of a protocol suite for quantum networks is still at its infancy [CacIllCal-23], and thus the functionalities of “quantum” upper-layers are yet to be defined.
@@
Fortunately, the communication limitations induced by the physical QLAN topology can be overcome, by relying on quantum entanglement, which enables a fundamentally new form of connectivity, referred to as entanglement-based connectivity [IllCalMan-22, MazCalCac-24-QCNC, MazCalCac-25]. Specifically, by leveraging multipartite entanglement, it is possible to establish a virtual topology – referred to as artificial topology – overlaying the physical one. The artificial topologies activated by entanglement can significantly differ from the underlying physical topology, and they can – if properly engineered – effectively mitigate the constraints imposed by the QLAN architecture.
@@
Indeed, as depicted in Fig. 6, the required artificial topology can be engineered to support a wide range of quantum communication patterns. However, the choice and the design of the multipartite entangled state to be generated and distributed is highly non-trivial and has great influence on the enabled network functionalities [IllCalVis-23, IllCalVis-23, CheIllCac-24]. As instance, it is well known that by sharing a GHZ state, a single Bell state can be extracted, on-demand, whenever needed, between any pair of network nodes sharing part of the entangled state [IllCalVis-23]. This has the advantage to dynamically adapt to the communication demands, without predetermining the identities of the involved nodes [^4] Conversely, when a Bell state is shared between two parties, the identities of nodes involved in the communication has been fixed a-priori.. Although GHZ states present interesting properties compared to the use of bipartite entangled states, they are limited by the extraction of a single Bell state, even if the number of qubits of the overall state grows arbitrarily. At the same time, GHZ states are characterized by unitary persistency [IllCacMan-21], i.e., the minimum number of qubits that need to be measured to guarantee that the resulting state is separable is one. Thus, they are highly sensitive to errors: losses of even one of the particles causes the entire state to lose coherence.
@@
Accordingly, the choice of the multipartite state to be distributed in the network involves fundamental compromises between three main factors: the number of Bell pairs that can be extracted – after manipulations of the initial state – persistency and robustness to quantum noise. Thus the degrees of freedom for the design can be resumed as follows [MazCalCac-25]:
@@
[…]
@@ 3.2.1 Graph states as network resources
[…]
@@
Figure 7: Schematic representation of the correspondence between graph states measurements and the associated graph operations (a). Application of Pauli-Y measurements for the engineering of the artificial topology of the network in a 66 clients QLAN topology (b).
@@ 3.2.2 From physical to artificial topologies
Accordingly to the above, graph states emerge as promising resource for engineering the QLAN topologies, due to their unique entanglement properties and operational flexibility[MazCalCac-24-QCNC, MazZhaChu-24, MazCalCac-25]. The centralized QLAN model, depicted in Fig. 6, allows the orchestrator node to generate and distribute graph states and linear graph states, while minimizing client-side operations and associated communication overhead. To this aim, it retains for itself part of the qubits of the generated graph state while distributing the remaining ones to the client nodes. With this strategy, the orchestrator node can engineer the artificial topology solely through the use of its locally held qubits and by employing Local Operations and Classical Communication (LOCC). Thus, this approach allows the dynamic reconfiguration of the QLAN topology, without requiring distributed quantum operations at the clients, besides the correction unitaries after the Pauli measurements, as indicated in Tab 2.
@@
The orchestrator can leverage the inherent simplicity of linear graph states as building block for synthesizing more sophisticated entangled states, as recently shown in [MazCalCac-25]. This approach enables the engineering of denser QLAN artificial topologies. This, in turn, implies that the orchestrator gains enhanced capabilities to tailor the connectivity patterns, allowing for more versatile and resource-efficient quantum computation architectures.
An example of dense resulting artificial topology is represented in Fig. 6(b). Remarkably, the presence of multiple entangled links in the artificial topology does not necessarily imply the extraction of more Bell states [MazCalCac-25], but rather simplifies the selection of the identities of the target nodes for dedicated link extractions. As a consequence, the orchestrator node is capable of enhancing the reliability and adaptability of the overlying entangled network topology, thanks to the possibility of reconfiguring paths according to the traffic demands – requests of specific entangled links – or node disconnections. Interestingly, as also represented in Fig. 8(a), the orchestrator is able to generate and distribute wisely manipulated linear graph states through merging [^5] The merging operation between vertices belonging to different graph states combines two vertices into a single one. This operation is equivalent to the so-called fusion when referred to optical graph state operations. operations. Specifically, the orchestrator is able to further engineer the resulting connectivity of the network nodes according to proper Pauli-X measurements. As depicted in Fig. 8(b), the adjacency between any two target clients can be achieved by reducing the distance of the clients in the resulting artificial topology. As a result, proximity engineering of physically unconnected clients has the direct consequence of simplifying the extraction of the dedicated entangled links in the artificial topology.